# VPRTempo: A Fast Temporally Encoded Spiking Neural Network for Visual Place Recognition

### v1

@@ III-F Evaluation metrics
III-G Baseline methods
We evaluated our system against standard VPR methods and state-of-the-art SNN networks. Firstly, the simple sum of absolute differences (SAD) baseline calculates the pixel-wise absolute difference between query and reference database images and selects a match based on the lowest sum value [36]. For SAD, images were resized to the same width and height used in our system (28 x 28 pixels). NetVLAD is trained to be highly viewpoint robust, and in this case images were kept at the original resolution [5]. We also compare with the current state-of-the-art Generalized Contrastive Loss (GCL) [30], which for Oxford RobotCar the dataset was resized to 240x320. We note that both Nordland and Oxford RobotCar are “on-the-rails” datasets with very limited viewpoint shift, which disadvantages GCL and NetVLAD. Finally, the main comparison is VPRSNN which is a recently reported SNN for performing VPR tasks based on spike rate coding by Hussaini et al. [21].
@@ III-G Baseline methods
Nordland (2700 places) Oxford RobotCar (360 places)
Method P@100R (%) Train C (min) Query C (Hz) Train. G (min) Query G (Hz) P@100R (%) Train. C (min) Query C (Hz) Train. G (min) Query G (Hz)
@@ III-H Hyperparameter search
To tune hyperparameters, we initially used a random search to identify the most influential parameters on match accuracy. Specifically, the hyperparameters we optimized were: firing threshold $\theta$, initial STDP learning rate $\eta_{\text{STDP}}^{\text{init}}$, initial ITP learning rate $\eta_{\text{ITP}}^{\text{init}}$, firing rate range $f_{min}$ and $f_{max}$, excitatory and inhibitory connection probabilities $P_{exc}$ and $P_{inh}$, and the constant input $C$. An initial random sweep of 5,000 identified parameters for a refined grid search, specifically for $f_{min}$, $f_{max}$, $P_{exc}$, and $P_{inh}$. The resulting hyperparameters can be found in Table I.
IV Results
@@
In this section, we establish the network training and query speeds when run on either CPU and GPU hardware (Section IV-A). Then, we compare the performance of the network when compared to state-of-the-art and standard VPR systems (VPRSNN [21], GCL [30], NetVLAD [5], and SAD [36]) (Section IV-B).
IV-A Training and querying speeds
To establish how our network performed against the previous state-of-the-art SNN for VPR (VPRSNN), we measured the total time taken to train each network on 3300 places from the Nordland dataset [21]. We also tested query speeds of both networks, and the effect the number of places has on training and querying for our network (Figure 2).
Figure 2: A Comparison of training times for 3300 places from the Nordland dataset our system against state-of-the-art (VPRSNN [21]). VPRSNN trained in 360 mins, VPRTempo on a CPU trained in 60 mins, with our best performance of $\approx$1 min when VPRTempo runs on a GPU. B Querying speed 2700 places: VPRSNN 2 Hz, VPRTempo CPU at 353 Hz, and VPRTempo GPU queried at 1634 Hz. C Increasing the number of places scales the training time with a time complexity of $\mathcal{O}(n)$ and inference time increases with a time complexity of $\mathcal{O}(log\ n)$.
@@ IV-A Training and querying speeds
VPRSNN learned 3300 places in approximately 360 mins compared with our system which only took 10% (60 mins) and 0.4% (1 mins) of the time to train an equivalent number of images when run on CPU and GPU hardware, respectively (Figure 2A). Whilst our network trained the fastest on GPU hardware, CPU training was still significantly faster than VPRSNN [21] indicating our system’s capability for real-time operation when training new places online. We attribute this predominately to increased spike efficiency and the shift away from modeling biological complexity, as in the Brian2 simulator [53].
@@
Our system is capable of real-time query deployment, with speeds on both CPU and GPU hardware for our network in the order of hundreds to thousands of Hz (Figure 2B), which is particularly useful for resource limited compute scenarios. A single query simultaneously activating all expert modules at these speeds shows the scalability of our method if larger datasets are employed. And finally, the time scaling of our system was found to be $\mathcal{O}(n)$ for training time and $\mathcal{O}(log\ n)$ for query speed (Figure 2C). We next measured training and query speeds for the conventional SAD and NetVLAD and state-of-the-art GCL VPR methods, as summarized in Table II. Across the board, our network trained and queried substantially faster than all of the other methods, with $\approx$ sub-hour training times on both CPU and GPU hardware and query speeds in the hundreds to thousands of hertz. This indicates that our network is suitable for real-time compute scenarios in resource limited systems.
We note that SAD has no training requirement to run and NetVLAD and GCL uses pre-trained models [5, 30]. Query speed for NetVLAD and GCL was measured as the time taken for query image feature extraction [5, 30]. Only CPU training or query speeds are listed for VPRSNN [21] since there is no GPU implementation of this network available.
IV-B Network accuracy
@@
Now that we have established the beneficial training and inference speed, we next evaluated our system accuracy when compared to conventional and state-of-the-art VPR systems. For the Nordland traversal dataset, our system achieved a P@100R of 56% (Table II). The conventional methods SAD and NetVLAD achieved an P@100R of 48% and 31% respectively, with the state-of-the-art GCL being 32%. Our main comparison competitor to VPRSNN achieving 53% (Table II). By comparison, for Oxford RobotCar dataset we achieved a P@100R of 37% compared to 40% for VPRSNN (best performer) (Table II). SAD, NetVLAD, and GCL measured at 37%, 31%, and 33% respectively. As previously discussed (refer to Section III-G), NetVLAD and GCL do not perform as well on static view-point datasets like Nordland and Oxford RobotCar. In addition to this, we calculated the precision-recall curves and R@N for both datasets to further validate our method, which are shown in Figure 3C and D.
@@
Figure 3: A Example database and query image from the Nordland dataset (left) that are patch-normalized for network training (right). B Ground truth (GT, left) and descriptor similarity matrix from testing 2700 query images (right). C Precision-recall curves comparing our network with sum of absolute differences (SAD [36]), NetVLAD [5], Generalized Contrastive Loss (GCL [30]), and VPRSNN [21]. D Recall at N curves comparing methods as in C.
V Conclusions and future work
@@
In this paper, we presented the first temporally encoded SNN to solve VPR tasks for robotic vision and localization. We significantly improved the capacity of SNN systems to learn and query place datasets over state-of-the-art and conventional systems (SAD [36], NetVLAD [5], GCL [30], VPRSNN [21]), demonstrating capability for real-time learning in deployment. In addition to this, we observed comparable network precision to these systems in correctly matching query to database references.

### v2

@@ IV-C Evaluation metrics
IV-D Baseline methods
We evaluated our system against standard VPR methods and state-of-the-art SNN networks. Sum of absolute differences (SAD) calculates the pixel-wise absolute difference between query and reference database images, selecting a match based on the lowest sum [33]. For SAD, images were resized to 28 x 28 pixels. NetVLAD is trained to be highly viewpoint robust, and in this case images were kept at the original resolution [11]. We also compare with the current state-of-the-art Generalized Contrastive Loss (GCL) [14]. For Oxford RobotCar the dataset was resized to 320x240. Nordland and Oxford RobotCar are “on-the-rails” datasets with very limited viewpoint shift, which disadvantages GCL and NetVLAD. Our main comparison is VPRSNN, a recently reported SNN for performing VPR tasks based on spike rate coding by Hussaini et al. [8].
@@ IV-D Baseline methods
Nordland (3300 places) Oxford RobotCar (450 places)
Method $R@1$ (%) Train CPU (min) Query CPU (Hz) Train GPU (min) Query GPU (Hz) $R@1$ (%) Train CPU (min) Query CPU (Hz) Train GPU (min) Query GPU (Hz)
@@ IV-E Hyperparameter search
To tune hyperparameters, we initially used a random search to identify the most influential parameters on match accuracy. Specifically, the hyperparameters we optimized were: firing threshold $\theta$, initial STDP learning rate $\eta_{\text{STDP}}^{\text{init}}$, initial ITP learning rate $\eta_{\text{ITP}}^{\text{init}}$, firing rate range $f_{min}\leq f_{j}^{n}\leq f_{max}$, excitatory and inhibitory connection probabilities $P_{exc}$ and $P_{inh}$, and the constant input $C$. An initial random sweep of 5,000 identified parameters for a refined grid search, specifically for $f_{min}$, $f_{max}$, $P_{exc}$, and $P_{inh}$. The resulting hyperparameters can be found in Table I and were used for both datasets. Hyperparameters were tuned on image subsets not used for any of the presented training and query performance results.
V Results
@@
In this section, we establish the network training and query speeds when run on either CPU and GPU hardware (Section V-A). Then, we compare the performance of the network when compared to state-of-the-art and standard VPR systems (VPRSNN [8], GCL [14], NetVLAD [11], and SAD [33]) (Section V-B).
V-A Training and querying speeds
@@
To establish network performance against a previous state-of-the-art SNN for VPR (VPRSNN), we measured the time to train each network on 3300 places from Nordland [8]. We also tested query speeds of both networks, and the effect the number of places has on both (Figure 2). VPRSNN learned 3300 places in approximately 360 mins, compared to our system which required 17% (60 mins) and 0.28% (1 mins) of the time to train an equivalent number of images when run on CPU and GPU hardware, an Intel i7-9700K and Nvidia RTX 2080 respectively (Figure 2A). Whilst our network trained the fastest on GPU hardware, CPU training was still significantly faster than VPRSNN [8]. We attribute this predominately to increased spike efficiency, shifting away from modeling biological complexity in SNN simulators [66].
Figure 2: A Comparison of training times for 3300 places from the Nordland dataset our system against state-of-the-art (VPRSNN [8]). VPRSNN trained in 360 mins, VPRTempo on a CPU trained in 60 mins, with our best performance of $\approx$1 min when VPRTempo runs on a GPU. B Querying speed 2700 places: VPRSNN 2 Hz, VPRTempo CPU at 353 Hz, and VPRTempo GPU queried at 1634 Hz. C Increasing the number of places scales the training time with a time complexity of $\mathcal{O}(n)$ and inference time increases with a time complexity of $\mathcal{O}(\log\!n)$.
@@
Capable of real-time query deployment, our system inferences on CPU and GPU hardware in the order of hundreds to thousands of Hz (Figure 2B), critical for resource limited compute scenarios. The time scaling of our system when training on reference datasets from 100 to 27,000 places was found to be $\mathcal{O}(n)$ for training and $\mathcal{O}(\log\!n)$ for query times, respectively (Figure 2C). We next measured training and query speeds for the conventional SAD and NetVLAD and state-of-the-art GCL VPR methods, as summarized in Table II. VPRTempo trained and queried substantially faster than all of the other methods. We note that SAD has no training requirement to run and NetVLAD and GCL uses pre-trained models [11, 14]. Query speed for NetVLAD and GCL was measured as the time taken for query image feature extraction [11, 14]. Only CPU training or query speeds are listed for VPRSNN [8] since there is no GPU implementation of this network available.
V-B Network accuracy
@@
Now that we have established the beneficial training and inference speed, we next evaluated our system accuracy when compared to conventional and state-of-the-art VPR systems. For the Nordland dataset, our system achieved a $R@1$ of 56% (Table II). The conventional methods SAD and NetVLAD achieved an $R@1$ of 48% and 31% respectively, with the state-of-the-art GCL being 32%. Our main comparison competitor to VPRSNN achieved 53% (Table II).
By comparison, for Oxford RobotCar dataset we achieved a $R@1$ of 37% compared to 40% for VPRSNN (best performer) (Table II). SAD, NetVLAD, and GCL measured at 37%, 31%, and 33% respectively. In addition to this, we calculated the precision-recall curves and $R@N$ for both datasets to further validate our method, which are shown in Figure 3C and D. We speculate that VPRTempo performs better in Nordland than Oxford RobotCar due to the lower viewpoint variation. As previously discussed (refer to Section IV-D), NetVLAD and GCL do not perform as well on static view-point datasets like Nordland and Oxford RobotCar.
@@
Figure 3: A Example database and query image from the Nordland dataset (left) that are patch-normalized for network training (right). B Ground truth (GT, left) and descriptor similarity matrix from testing 3300 query images (right). C Precision-recall curves comparing our network with sum of absolute differences (SAD [33]), NetVLAD [11], Generalized Contrastive Loss (GCL [14]), and VPRSNN [8]. D Recall at N curves comparing methods as in C.
VI Conclusions and future work
@@
In this paper, we presented the first temporally encoded SNN to solve VPR tasks for robotic localization and navigation. We significantly improved the capacity of SNN systems to learn and query place datasets over state-of-the-art and conventional systems (SAD [33], NetVLAD [11], GCL [14], VPRSNN [8]). In addition to this, we observed comparable network precision to these systems in correctly matching query to database references.